# Strong Coupling Constant Determination from the new CTEQ-TEA Global QCD Analysis

### v1

@@ Acknowledgments
The work of MG is partially supported by the National Science Foundation under Grants No. PHY-2412071. KM was supported in part by the US National Science Foundation under Grant PHY-2310497. PMN and SD are grateful for support from the Wu-Ki Tung Endowed Chair in particle physics. The work of SD was also supported by the National Natural Science Foundation of China under Grant number 12475079. The work of CPY is supported by the U.S. National Science Foundation under Grants No. PHY-2310291.
@@ Appendix A New LHC data sets included in the CT25 NNLO analysis
Lepton pair production measurements. As in the CT18A global fit, we include the precision measurement of $W$ and $Z$ production by ATLAS collaboration at 7 TeV [43]. This measurement revealed the enhancement in the strangeness PDF at $x\sim 0.02$ – the trend that was corroborated by more recent LHC observations. In Ref. [40], we examined in depth the implications of these measurements for PDFs and identified the following differential cross sections for inclusion in the final CT25 fit:
@@
      muon pseudorapidity distribution in $W$ boson production $(W\rightarrow\mu\nu)$ at a center-of-mass energy $\sqrt{s}=8$ TeV [26] with 20.2 fb-1 of integrated luminosity (IL), in the fiducial volume $p^{l,\nu}_{T}>25$ GeV, $|\eta_{\mu}|<2.4$. $m_{T}>40$ GeV;
@@
      $W^{\pm}$ muon pseudorapidity and $Z$ dilepton rapidity distributions [29] at $\sqrt{s}=5.02$ TeV with 25 pb-1 of IL, in the fiducial phase space $p^{l,\nu}_{T}>25$ GeV, $|\eta_{l}|<2.5$, $m_{T}>40$ GeV for $W^{\pm}$ bosons, and $p^{l}_{T}>20$ GeV, $|\eta_{l}|<2.5$, $66<m_{ll}<116$ GeV for $Z$ bosons;
@@
    From the CMS collaboration: rapidity distribution for the $Z$ boson production at 13 TeV [27] with an IL of 35.9 fb-1, $p^{l}_{T}>20$ GeV, $|\eta_{l}|<2.4$, and $|m_{ll}-M_{Z}|<15$ GeV.
@@
    From the CMS collaboration, we include the $d^{2}\sigma/dp_{T}d|y|$ distribution at $\sqrt{s}=13$ TeV [38] with 33.5 fb-1 of IL, R=0.7R={0.7}, $|y|<2$ and $97\leq p^{jet}_{T}\leq 3103$ GeV.
Appendix B Sensitivity to charm-quark mass variations
@@
To investigate the sensitivity of the results to the charm-quark mass $m_{c}$ (pole mass approximation), we extracted $\alpha_{s}(M_{Z})$ from two additional $\chi^{2}_{R}$ scans obtained with $m_{c}=1.2$ GeV and $m_{c}=1.4$ GeV respectively. The default charm-quark mass value used in the CT25 global analysis is $m_{c}=1.3$ GeV.

### v2

@@ Acknowledgments
We thank Amanda Cooper-Sarkar, Olaf Behnke, Tom Cridge, Louis Lyons, Robert Thorne, CTEQ-TEA colleagues, and participants of PDF4LHC-PHYSTAT meetings for conversations on uncertainty quantification for QCD parameters determined in PDF fits. We thank Tie-Jiun Hou for help on the CT25 analysis. The work of MG is partially supported by the National Science Foundation under Grant No. PHY-2412071. KM was supported in part by the US National Science Foundation under Grant PHY-2310497. PMN and SD are grateful for support from the Wu-Ki Tung Endowed Chair in particle physics. The work of SD was also supported by the National Natural Science Foundation of China under Grant number 12475079. The work of CPY is supported by the U.S. National Science Foundation under Grant No. PHY-2310291.
@@ Appendix A New LHC data sets included in the CT25 NNLO analysis
Lepton pair production measurements. As in the CT18A global fit, we include the precision measurement of $W$ and $Z$ production by ATLAS collaboration at 7 TeV [29]. This measurement revealed an enhancement in the strangeness PDF at $x\sim 0.02$ – a trend that was corroborated by more recent observations at the LHC. In Ref. [26], we examined in depth the implications of these measurements for PDFs and identified the following differential cross sections for inclusion in the final CT25 fit:
@@
      muon pseudorapidity distribution in $W$ boson production $(W\rightarrow\mu\nu)$ at a center-of-mass energy $\sqrt{s}=8$ TeV [40] with 20.2 fb-1 of integrated luminosity (IL), in fiducial volume $p^{l,\nu}_{T}>25$ GeV, $|\eta_{\mu}|<2.4$. $m_{T}>40$ GeV;
@@
      $W^{\pm}$ muon pseudorapidity and $Z$ dilepton rapidity distributions [43] in $\sqrt{s}=5.02$ TeV with 25 pb-1 of IL, in the fiducial phase space $p^{l,\nu}_{T}>25$ GeV, $|\eta_{l}|<2.5$, $m_{T}>40$ GeV for $W^{\pm}$ bosons and $p^{l}_{T}>20$ GeV, $|\eta_{l}|<2.5$, $66<m_{ll}<116$ GeV for $Z$ bosons;
@@
    From the CMS collaboration:
    rapidity distribution for $Z$ boson production at 13 TeV [41] with an IL of 35.9 fb-1, $p^{l}_{T}>20$ GeV, $|\eta_{l}|<2.4$ and $|m_{ll}-M_{Z}|<15$ GeV.
@@
    From the CMS collaboration,
    1. (a)
      In the CT25prel fit, we include $d^{2}\sigma/dp_{T}d|y|$ distributions at $\sqrt{s}=13$ TeV [52] with 33.5 fb-1 of IL, R=0.7R={0.7}, $|y|<2$ and $97\leq p^{jet}_{T}\leq 3103$ GeV.
    2. (b)
      In the CT25 fit, we include $d^{2}\sigma/(dp_{T}dy)$ distributions at $\sqrt{s}=7~\mathrm{TeV}$ with an IL of $5~\mathrm{fb}^{-1}$ (ID=556) [49, 50, 51, 6], at $\sqrt{s}=8~\mathrm{TeV}$ with an IL of $19.7~\mathrm{fb}^{-1}$ (ID=557) [52, 6], and at $\sqrt{s}=13~\mathrm{TeV}$ with an IL of 33.5 fb-1 (ID=555) [48, 6]. The $7~\mathrm{TeV}$ measurements include 118 data points in six rapidity bins covering $0\leq|y|\leq 2.5$, with jet transverse momenta in the range $114\leq p_{T}\leq 1327~\mathrm{GeV}$. The $8~\mathrm{TeV}$ measurements include 164 data points in six rapidity bins covering $0\leq|y|\leq 3.0$, with $74\leq p_{T}\leq 1784~\mathrm{GeV}$. The $13~\mathrm{TeV}$ measurements include 78 data points, $|y|<2$ and $97\leq p^{jet}_{T}\leq 3103$ GeV. These CMS 7, 8, 13 TeV measurements have updated systematic uncertainties and newly provided statistical correlation matrices [51, 6]. They supersede the earlier inclusive-jet datasets at $7~\mathrm{TeV}$ (ID=542) [49, 50], $8~\mathrm{TeV}$ (ID=545) [52], and $13~\mathrm{TeV}$ (ID=555) [48] included in previous analyses [16, 28].
Appendix B Investigation of reclustering prescriptions
In continuation of the discussion in Sec. IV.6.1, we illustrate the possibility of separate reclustering of the residuals and nuisance parameters by two figures. Figure 12 goes back to the example of the GT and DT uncertainties for the published data sets in Fig. 6 for the CT25 baseline and now shows the DT uncertainties constructed only with $D^{2}$ and only with $R^{2}$ for the CT25prel baseline. [^9] These uncertainties do not change much between the two baselines. In this and following figure, the lower bands saturate at $\alpha_{s}(M_{Z})=0.113$, the lowest point on our $\alpha_{s}$ grid. The GT and DT uncertainties based only on residuals (i.e., without $R^{2}$) are wider than the corresponding ones based on both residuals and nuisance parameters ($D^{2}+R^{2}$). The role of $R^{2}$ appears to be somewhat different for the GT and DT cases. In the DT case, the $R^{2}$ has a more constraining power on the uncertainty: if it were to be the only metric for the uncertainty, we would obtain a very small error (blue dashed band). However, $R^{2}$ is in general not perfectly determined, due to lack of full control on the correlated systematic uncertainties. The $R^{2}$ effect is less pronounced for the GT.
Figure 8 showed how the 68% and 90% CL error estimates change when we vary the DT clustering procesure. Figure 13 extends this comparison to show error estimates for other credibility levels, including the residuals only (that is, neglecting $R^{2}$). In this figure, the “Global T: no $R^{2}$” and “Dynamical T: no $R^{2}$” refer to the GT and DT uncertainties based on the full baseline. The “52 alt. clusters” and other lines refer to randomly generated clusters constructed by reshuffling the residuals in the 52 experimental data sets of CT25prel. For the alternative clusters, the lines show the envelopes of the DT estimates for 200 shuffles of the corresponding alternative clusters. In an ideal scenario, in the absence of tensions among the data, alternative cluster configurations would produce approximately the same estimate for the $\alpha_{s}(M_{Z})$ uncertainty. In reality, when we partition into a lot of clusters, taking the outcome of the DT without $R^{2}$ by itself may already produce a stringent (likely, too stringent) uncertainty on $\alpha_{s}(M_{Z})$. However, the constraints may be weaker for other cluster configurations, and including $R^{2}$ into clustering may additionally reduce or increase the strength of the DT constraints by modifying the tensions, similarly to what we saw for the DT that reclustered the published data sets.
Appendix C A toy example of uncertainty under K-folding
cluster 0 cluster 1 cluster 2 cluster 3
$N_{\textrm{pt}}$ 1113 1113 1112 1112
    $\chi^{2}_{R}$ (no $R^{2}$)
configuration 1 1094 1157 1281 1177
configuration 2 1102 1158 1252 1198
configuration 3 1065 1239 1251 1156
Table 4: Numbers of points and $\chi^{2}_{R}$ (without the systematic contribution $R^{2}$) for three random partitions (configurations) of the residuals into four pseudodata clusters.
This appendix provides details about the toy study of K-folding summarized in Sec. IV.6.2. Working with the best-fit residuals and nuisance parameters (assumed to be mutually independent) in the setup of Sec. IV.6, we divide the baseline data into four clusters according to randomly generated partitions (“configurations”). In contrast to the full K-folding procedure, we will not refit (optimize) the PDFs for each configuration, which would be too computationally expensive. Instead, we examine uncertainties by reclustering the output of the CT25prel global fit, similarly to what was done in the previous sections. In addition, to assess their independent impact, we will separately consider the residual and nuisance-parameter contributions.
Dependence of the central $\alpha_{s}(M_{Z})$ on clustering. We generate multiple partitions of the residuals from $\mathbf{v}_{r}$ into four clusters, with about 1113 points in each cluster, as outlined in Sec. IV.6.1. Table 4 reports the numbers of points and $\chi^{2}$ contributions for three representative configurations in our series. Similarly, we construct four $R^{2}$” clusters consisting of the shuffled best-fit nuisance parameters from $\mathbf{v}_{\lambda}$. Figure 9 illustrates the outcome of one such exercise, in which 100 configurations of four clusters were generated. The figure separately depicts the individual $\chi^{2}/N_{pt}$ ratios for the residuals (purple lines) and nuisance parameters (red-dashed lines). Each purple (red) line corresponds to a different configuration. The black lines represent the $\chi^{2}/N_{pt}$ average values. The stars represent the minima of the average curves which indicate the $\alpha_{s}(M_{Z})$ central-value preference. No tolerance criterion is applied here; we simply calculate the $\chi^{2}/N_{pt}$ for each configuration at different $\alpha_{s}(M_{Z})$.
configuration 1
in 1-clusters: {0} {1} {3} {2}
& 3-clusters: {0,1,2}, {0,1,3}, {0,2,3} {1,2,3} N/A N/A
configuration 2
in 1-clusters: {0} {1} {3} {2}
& 3-clusters: {0,1,2}, {0,1,3}, {0,2,3} {1,2,3} N/A N/A
configuration 3
in 1-clusters: {0} {3} {1} {2}
& 3-clusters: {0,1,2}, {0,1,3}, {0,2,3} {1,2,3} N/A N/A
Table 5: For each configuration of the K-folding exercise, we report the $\delta\alpha_{s}$ values found by applying the DT, as well as the one 1-cluster and three 3-clusters for which these $\delta\alpha_{s}$ are achieved. The columns containing the $\delta\alpha_{s}$ values and clusters are arranged so that the leftmost (rightmost) column contains the lowest (highest) $\delta\alpha_{s}$.
Just as one possibility, if the DT criterion is applied either to the “training set” consisting of 3 clusters or 1 “testing” cluster, it can render the uncertainty that can be substantially larger or smaller than the $^{+0.0007}_{-0.0014}$ from the NNPDF study. In its most granular realization, the DT criterion treats each residual as an independent “experiment”. Table 5 summarizes the outcome of applying such DT criterion to the three training clusters of the residuals (no $R^{2}$) and to the one left out. Here we show the DT $\alpha_{s}$ uncertainties for the same three configurations as in Table 4. These uncertainties can be either substantially smaller or larger than the NNPDF one, depending on the specific partition. The table also highlights the observation that, within a combination of any three clusters, the DT uncertainty of the whole cluster is equal to the smallest individual DT uncertainty among the constituent clusters. For this specific example, cluster 0 has the lowest $\chi^{2}$ for the three configurations, cf. Table 4. When combined with the other two, this cluster drives the net uncertainty of the combination. For the fourth 3-cluster combination, the uncertainty is set by the second most precise individual cluster. According to the findings in Table 5, the uncertainty on $\alpha_{s}(M_{Z})$ is not uniformly reproduced by each test cluster, and this is mainly ascribed to the tensions in the data. The results do not change if we consider the average of the errors over a large number of configurations.
Appendix D Information Criteria and the GMM
Figure 14: Variation of the AIC score with the error scaling factor $e_{S}$ shown for both $K=1$ (red) and $K=2$ (blue) in the GMM. Also shown are the values of $e_{S_{PDG}}\approx 4.7$ (dot-dashed) and $e_{S_{GMM}}\approx 6.2$ (dotted) – see the discussion in Sec. V.2.
We use the Akaike Information Criterion (AIC) to determine the optimal number of Gaussian mixtures $K$ for the GMM. The AIC score is defined as
$\text{AIC}=N_{p}\ \ln N_{\textrm{pt}}-2\ln L(\bar{\alpha}_{s})$ (30)
where $L(\bar{\alpha}_{s})$ is the GMM likelihood given in Eq. (26) evaluated at the optimized value of $\alpha_{s}$, $N_{\textrm{pt}}$ is the number of data clusters, and the number of independent parameters $N_{p}$ for the GMM is given by
$N_{p}=K\times N_{\text{par}}+(K-1)\ .$ (31)
Since we profile the $\chi^{2}$ over $\alpha_{s}(M_{Z})$, $N_{\text{par}}=1$. We vary the value of $K$ from $1$ through $5$ and find that the value of $K$ that minimizes the AIC score is $K=2$. Our GMM result in Eq. (27) quotes the $\alpha_{s}$ value for $K=2$.
This result is obtained for the actual input errors corresponding to error scaling factor $e_{S}=1$. In Fig. 14, we further illustrate how inflating the input uncertainties could modify the AIC analysis by plotting the AIC score vs. $e_{S}$ for two models: the usual $\sum_{i}\chi^{2}(\overline{\alpha}_{s},\alpha_{s,i},\sigma_{i})$ in Sec. IV.2, which corresponds to the choice $K=1$, and the GMM with $K=2$. The $K=2$ case achieves the lowest overall AIC score at $e_{S}\approx 1.2$, close to the nominal $e_{S}=1$.
For $e_{S}=1$, the $K=2$ model returns a weighted average of two distinct values of $\alpha_{s}(M_{Z})=\left\{{0.1154,0.1193}\right\}$. As $e_{S}$ increases, the data become more consistent, thus increasing the AIC score for $K=2$ and reducing it for $K=1$. Beyond the value of $e_{S}\approx 2.6$, the $K=1$ model would be a preferred choice. For $e_{S}\approx 2.6$, the fit is still of poor quality with a large $\sum_{i}\chi^{2}(\overline{\alpha}_{s},\alpha_{s,i},\sigma_{i})$, hence we do not use this as a measure to set the value of $e_{S}$ for $K=1$. For values of $e_{S}\gtrsim 3.5$, the AIC curves for both models grow linearly with the same slope. For these values of $e_{S}$, the $K=2$ model does not find two numerically distinct values of $\alpha_{s}(M_{Z})$ and therefore behaves like the $K=1$ model, albeit with some overfitting due to its larger number of parameters. This also explains a dip in the uncertainty for $K=2$ at $e_{S}\approx 4.2$ in Fig. 11, above which the $K=2$ uncertainty starts to track that for $K=1$. The slightly larger AIC score and uncertainties for $K=2$ for $e_{S}\gtrsim 3.5$ arise due to overfitting [86].
Appendix E Sensitivity to charm-quark mass variations
@@
To investigate the sensitivity of the results to the charm-quark pole mass $m_{c}$, we extracted $\alpha_{s}(M_{Z})$ from two additional $\chi^{2}_{R}$ scans obtained with $m_{c}=1.2$ GeV and $m_{c}=1.4$ GeV respectively. The default charm-quark mass value used in the CT25 global analysis is $m_{c}=1.3$ GeV.